# Is Long Context All You Need? Leveraging LLM’s Extended Context for NL2SQL Experiments, Analysis and Benchmark Paper

### v1

@@ 4.6. Self-correction with entire schema and more column values
Table 5. BIRD dev Ex Acc (%) and extra accumulated tokens for correction. The eval baseline pipeline does not include self-correction (SC); disambiguation refers to providing exntensive sample column values as described in Section 3.2; filtered schema is providing filtered relevant columns and values to re-write
@@
SC + filtered schema 65.84 2335.9111this excludes the cost of running the LLM-based column selection, where multiple LLM calls are needed to extract relevant columns for given request 35475.69
Table 5 shows how LLM’s self-correction can help with SQL generation and its refinement. We compare the LLM-based self-correction (SC) strategy with the ones with more advanced techniques, disambiguation and filtered schema. disambiguation is the long-context enabled technique where an extended lists of sample column values are shown to the model upon detecting empty results from syntactically correct SQL; filtered schema uses the column selection results mapping relevant schema elements per user request, as prescribed in (Talaei et al., 2024). Both disambiguation and filtered schema can help LLMs correct for any invalid literal and column column references. One potential issue with disambiguation is that the extra column values can distract the model, especially in the case of false positive detection (e.g., the answer should be null). We empirically show that it helps more and the technique performs slightly better or at least as good as SC. While comparable in performance, the techniques vary in terms of the cost. The accumulated average number of prompt tokens used for correction increases significantly for disambiguation, as with the full schema and extra sample column values. In general. accurate column selection can uplift NL2SQL performance, and the filtered schema does not increase the context size during correction.
However, column selection process to prepare filtered schema comes with retrieval cost (see Appendix A.2 for more discussion). filtered schema is a great strategy, if accurate column selection process is available. In our final long-context pipeline, we use SC + disambiguation to avoid extra retrieval/selection step and do everything in-context. We report the mean tokens/request plus 1-S.D, since the variance is large and to show how much tokens we use for the majority ( 68%) of the requests.
@@ 4.7. In-context learning with SQLite documentation
We test if the model can actually learn and improve its generation quality with SQLite documentation. We downloaded the entire SQLite documentation from its official homepage (sql, 2024). The original documentation comprises 16 million tokens across 792 HTML files, each covering a distinct theme. We applied two strategies to split the entire documentation into chunks so that we can augment the prompts with relevant information for ICL. The coarse-grained chucking strategy split the documentation by HTML file, while the fine-grained chucking strategy further separate each file by sections which ends up with 4125 smaller chunks in total. Similar to example retrieval, we embed the natural language questions and document chunks into vectors with the Gecko text embedding model (Lee et al., 2024b) and employ nearest neighbor search to identify the most relevant document chunks to the given question. We could further compress the context via summarization, but decided not to, since it drops important details, like syntax rules and examples.
Table 6. BIRD dev execution accuracy with SQLite document chunks, measured with the baseline pipeline.
@@
Ex Acc (%) 61.84 60.43 61.86 61.54 61.15
Context size (tok/req) 7380.87 7610.66 10115.78 14244.70 17560.80
We show the execution accuracy and context size for document retrieval in table 6(b). The retrieval is challenging because the natural language question does not reveal the full structure and features of the corresponding SQL query; furthermore writing a correct SQL query requires a combination of concepts from multiple sections from the documentation. However, we believe that the less than ideal retrieval process is not the reason why the documentation adds little to no value to generation. It is actually the nature of the errors that the LLMs make, which are mostly semantic errors (e.g., executable queries returning semantically irrelevant results).
The Gemini-1.5-Pro model is already well-versed with the SQLite syntax and function details that are illustarted in the documentation. The SQLite documentation is likely already in the pretrain mixture. Furthermore, the model can self-correct itself for syntactic errors based on the error message and without any documentation or examples. A recent work (Talaei et al., 2024) based on GPT-4 also points out that there is no obvious errors based on simple syntax misuse (e.g., function names), but rather more subtle formatting (e.g., does not follow required date format) and semantic errors (output SQL is executable but incorrect w.r.t. the request). Reading the SQLite documentation would not help with such semantic errors – because the model is already producing executable queries that are semantically incorrect. While there is no accuracy gain, putting the documentation chunks increase the context size, thus the latency, significantly.
@@ 4.8. Long context and latency relationship
Table 7. BIRD dev Ex Acc (%) and average generation and verification latency using the baseline pipeline with gemini-1.5-pro and gemini-1.5-flash
@@
Single generation latency (sec) 1.0 0.8
Single verification latency (sec) 0.9 0.2
Figure 5. Single LLM request latency by different context sizes. Notice that axes are in log-scale, and there is a step-wise linear relationship between latency and context size. Both gemini-1.5-pro and gemini-1.5-flash models suffer from increased and high-variable latency per request beyond the context size of $>32k$ tokens. The fitted lines represent best linear-fit between context size and latency measures.
In this section, we look at the context size (tokens) and latency relationship. Fig. 5 illustrates that there is near-linear relationship (strong positive correlation $R^{2}$=92.6). It is interesting to note that the latency and the variance starts increasing significantly beyond context size $>$32k. The larger context LLMs require inference computation scattered across more host and accelerators, which incurs more queueing delay. While we expect the smaller gemini-1.5-flash variant to have lower latency, due to this queueing delay and resource allocation differences it also suffers from increased latency at the long tail. The generation latency increases significantly (>>much-greater-than$>$ > 4 seconds) with larger context size ($>$32k tokens), thus we should increase the context only if there is a clear gain in generation quality. The average measures between context size and latency are linearly correlated, which makes it easy to model.
Table 7 illustrates the average latency difference for the two gemini-1.5 model variances, large and more expensive pro and small and cheaper flash. The difference in verification latency is more pronounced because the context size is smaller, whereas the generation with larger context pro and flash requires more prefill computation, which results in increased and similar queueing delay. Both models experience increased and similar average generation latency. For the single verification latency (indicative of the smaller context latency) flash is almost 75% faster than pro.
@@ 5. Ablation Study
Table 8. Ablation analysis of the long-context NL2SQL pipeline. The pipeline shows strong performance on BIRD dev, even without any fine-tuning and/or multiple candidate generation (a.k.a., self-consistency). The context size per request also increases significantly, and we measure the total accumulated tokens over any refinement iterations (self-correction) per given user request. We report the 1-standard-deviation from the mean stats to bound the majority of the requests.
@@
Overall, we focus on identifying useful information for improving NL2SQL performance – and measure the impacts of the contextual information in terms of both performance and cost. We exclude the rules and SQLite documentation chunks from the ablation study, since their separate contributions are negligible. The rules are fixed and included as part of the instructions. We also skip popular techniques, like Chain-of-Thought (CoT) prompting and self-consistency. CoT can help LLMs by breaking down a complex task into sub-tasks enabling multi-step reasoning to solve complex reasoning task, like NL2SQL (Liu and Tan, 2023; Tai et al., 2023); however, we observe that CoT does not improve the final output quality while increasing the context size by a few thousands tokens.
@@
Table 8 reports the generation performance (Ex Acc) along with accumulated context size (tokens/req) and latency (sec/req) per user request. The study is done adding one technique or extra contextual information at a time. The techniques and information that are more commonly used for NL2SQL are added earlier (schema, hints, column samples and self-correction, respectively) and our long-context pipeline specific components are added later (disambiguation and synthetic examples).
@@
Because the latest long-context LLM, gemini-pro-1.5, has really strong retrieval capability, we see the overall performance increase with more added information. However, when we break down the questions from BIRD dev into sub-categories by difficulties, we also see that some information are more helpful for simpler or harder requests versus the others. For instance, providing extra sample column values (or providing them all for disambiguation) is more helpful for challenging questions, where the question is more nuanced and/or ambiguous that seeing exact column values can help pick the right column and join paths. Synthetic examples show the opposite where it helps with simple and moderate questions, but hurt the challenging questions. Synthetic examples are generated to illustrate common SQL features and clauses using the schema elements from the target DB (Pourreza et al., 2024); the examples tend to be simpler and have less ambiguous mappings from the question and the SQL as the examples questions are generated by the same model describing its generated SQL. The hints are also interesting in that it appears to be one of the most critical ingredients for accurate SQL generation and it benefits moderate questions the most. BIRD datasets contain hints as to clarify nuanced concepts (e.g., “eligible free rate for K-12”) with a correct expression (“Eligible free rate for K-12 = ‘Free Meal Count (K-12)‘ / ‘Enrollment (K-12)‘”), and moderate queries often consist of more nuanced concepts that require such clarifications.
@@
One key aspect of leveraging long context is the cost. Table 8 illustrates how each long context information or technique contributes to the overall context size as the latency, which are highly correlated (we discuss the near-linear relationship between the context size and the latency in Section 4.8). For both context size $L$ and latency $T$, we report the average plus 1-S.D. because the variance of the measures are quite high given the retries and self-correction mechanisms. This way, the reported measures capture the majority of the requests we evaluated. Note that $L$ and $T$ track accumulated tokens and latency per request, until the final output is returned. Self-correction (retry) can be triggered for various reasons, including simple syntax errors to empty results. The pipeline will retry up to 5 times with appropriate fix/modification and higher temperature. For such recursive nature, self-correction increases the latency (and the accumulated tokens) significantly; the cost grow more exponentially with disambiguation (oversized column value examples). The online synthetic example generation also adds to the runtime latency significantly, since the generation process involves a long sequence of auto-regressive decoding. If synthetic examples are generated offline and retrieved at runtime , as tested in Section 4.5, then the generation latency can be reduced. It is interesting to note that adding a few thousands tokens does not increase the latency that significantly over the entire pipeline. For instance, increase of 8816 tokens per request for “verify & retry” delayed the latency only by 1.3 normalized latency units (mean + 1-S.D.).
@@ 6.2. Further performance improvement via self-consistency
Table 9. Long-context NL2SQL BIRD dev Ex Acc upper bounds (at least one of the candidates is correct) via multiple output candidates generation.) guarantees perfect recall
@@
Self-consistency is another very popular technique for NL2SQL (Wang et al., 2023; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024), where multiple candidates are generated and the most consistent answer or the best one is selected by a fine-tuned picker. As evidenced by the latest state-of-the-arts (Pourreza et al., 2024; Gao et al., 2024) on the BIRD leaderboard using this technique, generating multiple candidates, with different generators (a.k.a. NL2SQL pipelines) and each generator trying multiple times, and choosing the best one has become a crucial technique for achieving high accuracy in NL2SQL. One of the state-of-the-arts, CHASE-SQL, in Table 1 uses our long context pipeline as one of the three candidate generators to yield one of the strongest results on the BIRD leaderboard. One caveat of self-consistency is that this can quickly become expensive in terms of latency and the number of LLM calls. Self-consistency is orthogonal to our work, and can further improve the NL2SQL performance if used together with our pipeline. Table 9 illustrates how repeatedly generating multiple output candidates from our long context pipeline can perform (generation temperature set to 0.5). With an oracle picker (i.e., at least one of the candidates is correct per question) we can uplift the accuracy beyond 70%. In (Pourreza et al., 2024; Gao et al., 2024) the candidate sets are generated using multiple pipelines as different pipelines with different strategies result in more diverse outputs, boosting the final accuracy of self-consistency variants even further.
Self-consistency is orthogonal to the use of long context and outside the scope of this paper. In this work we focus on the quality of individual output candidate and study the performance implications of leveraging long context (extra contextual information), which has not been explored for NL2SQL before.

### v3

@@ 4.6. Self-correction with entire schema and more column values
Table 5. Ex Acc (%) and extra accumulated tokens for correction. The eval baseline pipeline does not include self-correction (SC); disambiguation refers to providing extensive sample column values as described in Section 3.2; filtered schema is providing filtered relevant columns and values to re-write
@@
SC + filtered schema 65.84 3930.55a 7001.89
KaggleDBQA test
Baseline (w/o SC) 58.91 - -
SC 59.45 2211.34 34573.67
SC + disambiguation 61.08 111238.44 205852.86
SC + filtered schema 59.45 2335.9111this excludes the cost of running the LLM-based column selection, where multiple LLM calls are needed to extract relevant columns for given request 35475.69
Table 5 shows how LLM’s self-correction can help with SQL generation and its refinement. We compare LLM-based self-correction (SC) with two more advanced techniques: disambiguation and filtered schema. disambiguation is the long-context enabled technique where an extended lists of sample column values are shown to the model upon detecting empty results from syntactically correct SQL; filtered schema uses the column selection results to filter relevant schema elements per user question, following the approach of (Talaei et al., 2024) . Both disambiguation and filtered schema can help LLMs correct for any invalid literal and column references. One potential issue with disambiguation is that the extra column values can distract the model, especially in the case of false positive detection (e.g., when the correct answer is null). However, empirical results show that it generally improves performance, performing slightly better or at least as well as SC. While comparable in performance, the techniques vary in terms of the cost. The accumulated average number of prompt tokens used for correction increases significantly for disambiguation, as it includes full schema and extra sample column values.
Accurate column selection is crucial for improving NL2SQL performance, and filtered schema does not increase the context size during correction. However, preparing filtered schema requires an additional retrieval step, which incurs a cost (see Appendix A.2 for more discussion). filtered schema is a great strategy, if accurate column selection process is available. However, in our final long-context pipeline, we use SC + disambiguation to avoid extra retrieval/selection step and do everything in-context. Due to the high variance in token usage, we report the mean tokens per request plus one standard deviation (1-S.D) to capture the range used for most requests ( 68%).
@@ 4.7. In-context learning with SQLite documentation
We test whether the model can effectively learn from SQLite documentation and improve its SQL generation quality. We downloaded the entire SQLite documentation from its official homepage (sql, 2024). The original documentation comprises 16 million tokens across 792 HTML files, each covering a distinct theme. We applied two strategies to split the entire documentation into chunks so that we can augment the prompts with relevant information for ICL. The coarse-grained chucking strategy split the documentation by HTML file, while the fine-grained chucking strategy further separate each file by sections which ends up with 4125 smaller chunks in total. Similar to example retrieval, we embed the natural language questions and document chunks into vectors with the Gecko text embedding model (Lee et al., 2024b) and employ nearest neighbor search to identify the most relevant document chunks to the given question. We could further compress the context via summarization, but decided not to, since it drops important details, like syntax rules and examples.
Table 6. Ex Acc with SQLite document chunks, measured with the baseline pipeline.
@@
BIRD dev Ex Acc (%) 61.84 60.43 61.86 61.54 61.15
$\bar{L}$ (tok/req) 7380.87 7610.66 10115.78 14244.70 17560.80
KaggleDBQA test Ex Acc (%) 57.83 59.45 57.29 58.37 57.29
$\bar{L}$ (tok/req) 922.29 2187.10 6202.0 7240.81 9545.82
We show the execution accuracy and context size for document retrieval in table 6(b). The retrieval is challenging because the natural language question does not reveal the full structure and features of the corresponding SQL query; furthermore writing a correct SQL query requires a combination of concepts from multiple sections of the documentation. However, we believe that the imperfect retrieval is not the reason why the documentation adds little to no value to generation. It is actually the nature of the errors that the LLMs make, which are mostly semantic errors (e.g., executable queries returning semantically irrelevant results).
The Gemini-1.5-Pro model is already well-versed with the SQLite syntax and function details that are illustrated in the documentation. The SQLite documentation is likely already in the pre-train mixture. Furthermore, the model can self-correct itself for syntactic errors based on the error message and without any documentation or examples. A recent work (Talaei et al., 2024) based on GPT-4 also points out that there is no obvious errors based on simple syntax misuse (e.g., function names), but rather more subtle formatting (e.g., does not follow required date format) and semantic errors (output SQL is executable but incorrect w.r.t. the request). Reading the SQLite documentation would not help with such semantic errors – because the model is already producing executable queries that are semantically incorrect. While there is no accuracy gain, putting the documentation chunks increase the context size, thus the latency, significantly.
@@ 4.8. Long context and latency relationship
Table 7. Average normalized generation and verification latency on BIRD dev using the baseline pipeline with gemini-1.5-pro and gemini-1.5-flash. The normalized latency time unit is set to the average generation latency on BIRD dev.
@@
Single generation latency (units/req) 1.0 0.8
Single verification latency (units/req) 0.9 0.2
Figure 5. Single LLM request latency by different context sizes. The axes are in log-scale, and there is a step-wise linear relationship between latency and context size. Both pro and flash models suffer from increased and high-variance latency per request beyond $>32k$ tokens context size.
In this section, we look at the context size (tokens) and latency relationship. The relationship between context size and latency in Google’s gemini-1.5 generation API is primarily a function of the number of tokens and is agnostic to the benchmark datasets. We evaluate this relationship using sampled requests of varying context size from BIRD dev. Figure 5 illustrates that there is near-linear relationship (strong positive correlation $R^{2}$=92.6). It is interesting to note that the latency and the variance starts increasing significantly beyond context size $>$32k. The larger context LLMs require inference computation scattered across multiple hosts and accelerators, introducing additional queuing delays. While we expect the smaller gemini-1.5-flash variant to exhibit lower latency, it also suffers from increased latency in the long tail due to queuing delays and differences in resource allocation. The generation latency increases significantly (>>much-greater-than$>$ > 4 seconds) with larger context size ($>$32k tokens), thus we should increase the context only if there is a clear gain in generation quality. Fortunately, the average relationship between context size and latency remains linear, making it easy to model.
Table 7 illustrates the average latency difference for the two gemini-1.5 model variants: the larger, more expensive pro and the small, more cost-efficient flash. The average generation latency for BIRD dev is used as the unit for normalized latency $T$. The difference in verification latency is more pronounced because the context size is smaller, whereas the generation with larger context pro and flash requires more prefill computation, which results in increased and similar queuing delay. Both models experience increased and similar average generation latency. For the single verification latency (indicative of the smaller context latency) flash is almost 75% faster than pro.
@@ 5. Ablation Study
Table 8. Ablation analysis of the long-context NL2SQL pipeline. The pipeline shows strong performance on BIRD dev, even without any fine-tuning and/or multiple candidate generation (a.k.a., self-consistency). The context size per request also increases significantly, and we measure the total accumulated tokens over any refinement iterations (self-correction) per given user request. We report 1-standard-deviation from the mean stats to bound majority of the requests.
@@
Overall, we focus on identifying useful information for improving NL2SQL performance – and measure the impacts of the contextual information in terms of both performance and cost. We exclude the rules and SQLite documentation chunks from the ablation study, since their separate contributions are negligible. The rules are fixed and included as part of the instructions. We also omit certain popular techniques, such as Chain-of-Thought (CoT) prompting and self-consistency. CoT aids LLMs by breaking down complex tasks into subtasks, enabling multi-step reasoning for structured problems like NL2SQL (Liu and Tan, 2023; Tai et al., 2023). However, our observations indicate that CoT does not improve final output quality while significantly increasing context size—often by several thousand tokens.
@@
Table 8 reports the generation performance (Ex Acc) along with accumulated context size (tokens/req) and latency (sec/req) per user request using BIRD dev – the most widely used NL2SQL benchmark, containing diverse questions and tables spanning multiple domains. We also report the results of a mini-ablation study using other benchmark datasets in Table 9. The study follows an incremental approach, adding one technique or contextual information element at a time. The techniques and information that are more commonly used for NL2SQL are added earlier (schema, hints, column samples and self-correction, respectively) and our long-context pipeline specific components are added later (disambiguation and synthetic examples).
@@
Because the latest long-context LLM, gemini-pro-1.5, has strong retrieval capabilities, we observe overall performance improvements as more contextual information is included. However, when we categorize BIRD dev questions by difficulty level, we find that different types of contextual information benefit different types of queries. For instance, providing extra sample column values (or providing them all for disambiguation) is more helpful for challenging questions, where the question is more nuanced and/or ambiguous - allowing the model to leverage exact column values to choose the correct column and join paths. Synthetic examples show the opposite where it helps with simple and moderate questions, but hurt the challenging questions. Synthetic examples are generated to illustrate common SQL features and clauses using the schema elements from the target DB (Pourreza et al., 2024). They tend to be simpler and involve less ambiguous mappings between natural language questions and SQL queries. The hints are also interesting in that it appears to be one of the most critical ingredients for accurate SQL generation and it benefits moderate questions the most. BIRD datasets contain hints as to clarify nuanced concepts (e.g., “eligible free rate for K-12”) with a correct expression (“Eligible free rate for K-12 = ‘Free Meal Count (K-12)‘ / ‘Enrollment (K-12)‘”), and moderate queries often consist of more nuanced concepts that require such clarifications.
@@
One key aspect of leveraging long context is cost. Table 8 illustrates how each long context information or technique contributes to the overall context size and the latency, which are highly correlated (we discuss the near-linear relationship between context size and latency in Section 4.8). For both context size $L$ and latency $T$, we report the average plus 1-S.D. because variance is high due to retries and self-correction mechanisms. This way, the reported measures capture the majority of the evaluated requests. Note that $L$ and $T$ track accumulated tokens and latency per request, until the final output is returned. Self-correction (retry) can be triggered for various reasons, ranging from simple syntax errors to empty results. The pipeline retries up to 5 times, applying appropriate fixes/modifications and increasing the temperature. Due to its recursive nature, self-correction increases the latency (and the accumulated tokens) significantly; the cost grows even more exponentially with disambiguation (which involves including additional column value examples). The online synthetic example generation also adds to the runtime latency significantly, since the generation process involves a long sequence of auto-regressive decoding. However, if synthetic examples are generated offline and retrieved at runtime, as tested in Section 4.5, the generation latency can be reduced. It is interesting to note that adding a few thousands tokens does not increase the latency that significantly over the entire pipeline. For instance, increase of 8816 tokens per request for “verify & retry” delayed the latency only by 1.3 normalized latency units (mean + 1-S.D.).
Table 9. A mini-ablation study analyzing the efficacy of disambiguation, many-shot ICL with synthetic examples for Long Context NL2SQL. The differences (in parentheses) are measured from the full pipeline results in Table 2.
w/o disambiguation w/o synthetic examples
Ex Acc (%) $\bar{T_{g}}$ (T units/req) Ex Acc (%) $\bar{T_{g}}$ (T units/req)
SPIDER 1.0 test 87.1 ( - ) 1.3 ( $\downarrow$ 0.2) 86.3 ( $\downarrow$ 0.8) 1.4 ( $\downarrow$ 0.1)
KaggleDBQA test 60.5 ( $\downarrow$ 0.6) 2.5 ( $\downarrow$ 0.2) 60.5 ( $\downarrow$ 0.6) 2.3 ( $\downarrow$ 0.4)
BEAVER dw 52.1 ( $\downarrow$ 8.3) 22.5 ( $\downarrow$ 467.3) 60.4 ( -) 595.5 ( $\uparrow$ 25.5)
Table 10. Further breakdown of mini-ablation study results, Ex Acc (%). As shown in Table 8 disambiguation (w/o synthetic examples) is more helpful for challenging or extra hard questions, while many-shot ICL with synthetic examples (w/o disambiguation) is more helpful for easier questions.
Question difficulty Easy Medium Hard Extra Hard
SPIDER 1.0 test w/o disambiguation 67.56 61.62 64.86 43.24
w/o synthetic examples 65.94 57.83 64.86 50.81
KaggleDBQA test w/o disambiguation 92.09 89.59 87.00 74.19
w/o synthetic examples 92.79 86.79 86.79 75.89
Table 9 reports the mini-ablation study, which analyzes the efficacy of the most controversial (expensive) long context techniques - disambiguation and many-shot ICL with synthetic examples - across other benchmark datasets. Running the full pipeline without each of the aforementioned techniques reduces both accuracy and latency (or context size). The performance implications of skipping disambiguation (w/o disambiguation) are limited on SPIDER and KaggleDBQA benchmarks, but much more pronounced on the BEAVER dataset. This is because disambiguation is only triggered for executable SQL queries that return empty result sets. Such invalid predictions happens more frequently in the BEAVER data warehouse (dw) dataset, as its schema is significantly more complex and requires more joins (4.25 on the average) compared to other datasets. Similarly, skipping many-shot ICL with synthetic examples (w/o synthetic examples) reduces both accuracy and latency, but the impact is limited on SPIDER and KaggleDBQA. On the contrary, w/o synthetic examples on BEAVER does not change the accuracy, but also increasing the latency significantly due to more frequent disambiguation attempts. This suggests that, overall, the generated examples are more misleading than helpful, but resulted in fewer empty results. This is expected, as the synthetic example generation primarily targets simpler query structures (Pourreza et al., 2024) rather than star-schema (data warehouse) queries involving many joins. It is also interesting to note that disambiguation and synthetic examples have very different performance characteristics. As shown in the full ablation study with BIRD dev (Table 8), disambiguation is more helpful for challenging questions, whereas synthetic examples contribute more to solving simpler problems. We observe the same trends across the other benchmark datasets, as illustrated in Table 10.
@@ 6.2. Further performance improvement via self-consistency
Table 11. Long-context NL2SQL BIRD dev Ex Acc upper bounds (at least one of the candidates is correct) with multiple candidate generation.
@@
Self-consistency is another very popular technique for NL2SQL (Wang et al., 2023; Talaei et al., 2024; Pourreza et al., 2024; Dong et al., 2023; Gao et al., 2023; Gao et al., 2024), where multiple candidates are generated, and the most consistent answer or the best one is selected by a fine-tuned picker. As demonstrated by the recent state-of-the-art approaches (Pourreza et al., 2024; Gao et al., 2024) on the BIRD leaderboard, self-consistency has become a crucial technique for achieving high accuracy in NL2SQL. This involves generating multiple candidates using different generators (a.k.a. NL2SQL pipelines) and each generator trying multiple times, and choosing the best one. One of the state-of-the-arts, CHASE-SQL, in Table 1 uses our long context pipeline as one of the three candidate generators, contributing to its top-tier performance on the BIRD leaderboard. One caveat of self-consistency is that it can quickly become expensive in terms of latency and the number of LLM calls. Self-consistency is orthogonal to our work, and can further improve the NL2SQL performance if used together with our pipeline. Table 11 illustrates the impact of generating multiple output candidates from our long context pipeline using a generation temperature of 0.5. With an oracle picker (i.e., at least one of the candidates is correct per question) we can uplift the accuracy beyond 70%. In (Pourreza et al., 2024; Gao et al., 2024), the candidate sets are generated using multiple pipelines, as different strategies result in more diverse outputs, further boosting the final accuracy of self-consistency.
Self-consistency is orthogonal to the use of long context and falls outside the scope of this paper. In this work, we focus on generating one high quality candidate and study the performance implications of leveraging long context (i.e., extra contextual information), that has not been previously explored for NL2SQL.